# I Predict Therefore I Am: Is Next Token Prediction Enough to Learn Human-Interpretable Concepts from Data?

### v2

@@ Abstract
The remarkable achievements of large language models (LLMs) have led many to conclude that they exhibit a form of intelligence. This is as opposed to explanations of their capabilities based on their ability to perform relatively simple manipulations of vast volumes of data. To illuminate the distinction between these explanations, we introduce a novel generative model that generates tokens on the basis of human-interpretable concepts represented as latent discrete variables. Under mild conditions—even when the mapping from the latent space to the observed space is non-invertible—we establish an identifiability result: the representations learned by LLMs through next-token prediction can be approximately modeled as the logarithm of the posterior probabilities of these latent discrete concepts, up to an invertible linear transformation. This theoretical finding not only provides evidence that LLMs capture underlying generative factors, but also strongly reinforces the linear representation hypothesis, which posits that LLMs learn linear representations of human-interpretable concepts. Empirically, we validate our theoretical results through evaluations on both simulation data and the Pythia, Llama, and DeepSeek model families.
@@ 1 Introduction
Large language models (LLMs) are trained on extensive datasets, primarily sourced from the Internet, enabling them to excel in a wide range of downstream tasks, such as language translation, text summarization, and question answering (Zhao et al., 2024; Bommasani et al., 2021; Olah et al., 2020). This remarkable success has ignited a heated debate surrounding the question of whether their performance is primarily achieved through simple operations applied to vast amounts of memorized data, or is a sign of genuine “intelligence”. Intelligence in this context would imply that the LLM had learned a model of the underlying generative factors that gave rise the data.
Recent empirical evidence has made it increasingly clear that the representations learned by LLMs encapsulate latent concepts, such as sentiment (Turner et al., 2023) or writing style (Lyu et al., 2023), which align with human-interpretable abstractions (Acerbi & Stubbersfield, 2023; Manning et al., 2020; Sajjad et al., 2022). Despite this, LLMs are often viewed as mere memorization systems, with their intelligence questioned due to the fundamental simplicity of next-token prediction. Critics argue that this seemingly trivial approach cannot explain the emergence of advanced AI capabilities, much less form the foundation of artificial general intelligence (AGI).
@@
Figure 1: An overview of the main contributions of this work. On the left, we illustrate the proposed latent variable model that represents concepts as latent variables $\mathbf{c}$, which are used to generate both the input $\mathbf{x}$ and output $y$ within a next-token prediction framework. Leveraging Bayes’ rule, the next-token prediction framework, and the diversity condition, we establish an identifiability result. Specifically, we demonstrate that the representations learned by LLMs approximately correspond to a linear transformation of the logarithm of the posterior distribution of latent variables conditioned on input tokens, i.e., $\mathbf{f}_{\mathbf{x}}(\mathbf{x})=\mathbf{A}[\log{p(\mathbf{c}=\mathbf{c}_{i}|\mathbf{x})}]_{\mathbf{c}_{i}}+o({\epsilon})$, where $o(\epsilon)$ represents a term that grows asymptotically smaller than $\epsilon$ as $\epsilon\to 0$. This identifiability results provide strong support for the linear representation hypothesis.
@@
Empirical studies suggest that LLM representations capture human-interpretable concepts in latent space (Turner et al., 2023; Acerbi & Stubbersfield, 2023; Manning et al., 2020; Sajjad et al., 2022). To analyze whether and how LLMs learn this through next-token prediction, we introduce a latent variable model, as shown on the left in Figure 1. We model latent concepts as variables connected by arbitrary causal relationships that give rise to observed variables, capturing the underlying generative process between concepts and observations. Although efforts have been taken to model the concepts as continuous linear vectors (Rajendran et al., 2024) or explore the empirically recognized linear geometry corresponding to the concepts (Park et al., 2023), we aim to develop a model enabling flexibility in capturing the data generative process and theoretical identifiability analysis. We explicitly model all variables, both latent and observed, as discrete, aligning with the inherently discrete nature of language and human-interpretable semantic representations of the world with discreteness. Specifically, we do not impose invertibility on the mapping from latent to observed space, as real-world data often involves complex relationships that strict invertibility constraints cannot capture. Building on this model, we develop a theoretical framework to assess whether next-token prediction can learn interpretable concepts. Through identifiability analysis, we show that, under mild conditions, representations learned by next-token prediction approximately correspond to a linear transformation of the logarithm of the posterior distribution of latent variables conditioned on input tokens. This connection offers theoretical insights into LLM mechanisms, highlighting their potential to model aspects of human intelligence. Surprisingly, our identifiability analysis offers a unified perspective supporting the linear representation hypothesis, which empirically suggests that concepts in LLMs are often encoded linearly (Mikolov et al., 2013; Pennington et al., 2014; Arora et al., 2016; Elhage et al., 2022; Burns et al., 2022; Tigges et al., 2023; Nanda et al., 2023; Moschella et al., 2022; Park et al., 2023; Li et al., 2024).
Our main contributions are summarized as follows:
@@
  • In Section 5, we present experiments on both simulation data, and real data across the Pythia (Biderman et al., 2023), Llama (Touvron et al., 2023a; Dubey et al., 2024), and DeepSeek-R1 (Guo et al., 2025) models families, with results consistent with our findings.
@@ 2 A Latent Variable Modal for Text Data
$[\mathbf{x},y]=\mathbf{g}(\mathbf{c}),\quad\text{or probabilistically},\quad p(\mathbf{x},y)=\sum_{\mathbf{c}}p(\mathbf{x}|\mathbf{c})p(y|\mathbf{c})p(\mathbf{c}),$ (1)
where latent variables $\mathbf{c}$ model latent human-interpretable concepts, and observed variables $\mathbf{x}$ and $y$ model words generated by the deterministic mapping $\mathbf{g}$ on $\mathbf{c}$. We here distinguish observed variables into $\mathbf{x}$ and $y$ to align the input, i.e., context and output token in the next-token prediction inference framework, respectively.
At first glance, modeling latent concepts as latent variables might seem inconsistent with empirical findings suggesting that concepts in LLMs are often linearly encoded, as posited by the linear representation hypothesis (Mikolov et al., 2013; Pennington et al., 2014; Arora et al., 2016; Elhage et al., 2022; Burns et al., 2022; Tigges et al., 2023; Nanda et al., 2023; Moschella et al., 2022; Park et al., 2023; Li et al., 2024; Gurnee et al., 2023). However, as we will demonstrate in Section 4.1, this modeling approach not only aligns with the linear representation hypothesis but also provides theoretical support for it.
Our goal is to impose realistic assumptions on the proposed latent variable model, ensuring the model’s flexibility to closely approximate real language data. To this end, we do not impose any specific graph structure between the latent variables. For example, we allow for any directed acyclic graph structure in the relationships between latent variables from the perspective of causal representation learning. Most importantly, we emphasize the following factors.
@@
Existing works predominantly assume that both latent and observed variables for modeling language data are continuous (Rajendran et al., 2024; Yan et al., 2024), primarily to simplify identifiability analysis. However, this assumption overlooks the inherently discrete nature of text. In contrast, we assume throughout this work that all variables, both latent variables $\mathbf{c}$ and observed $\mathbf{x}$ and $y$, are discrete. This assumption is particularly natural for observed variables $\mathbf{x}$ and $y$, given the inherently discrete structure of words and categories in text. Likewise, latent variables, which represent underlying concepts, are inherently well-suited to discrete modeling, as they often correspond to categorical distinctions or finite sets of semantic properties that naturally arise in text data. For instance, in a topic modeling scenario, latent variables can represent distinct topics, such as “sports”, “politics”, or “technology”, each of which forms a discrete category. This discrete structure aligns with the way humans typically interpret and classify information in text, making the assumption not only intuitive but also practical for real-world applications (Blei et al., 2003; Jin et al., 2020; Jiang et al., 2024). A recent study is constrained by its exclusive focus on binary concepts (Jiang et al., 2024), whereas our model removes this limitation, enabling the modeling of more complex and diverse latent structures. For example, latent variables could represent multi-class categories like age groups, various sentiment intensities, or a broader range of topic clusters, capturing richer and more nuanced concepts.
@@
Unlike many existing approaches (Jiang et al., 2024; Rajendran et al., 2024), we do not require the mapping from latent to observed variables to be invertible. This includes that which maps $\mathbf{c}$ to generate words $\mathbf{x}$, and that which maps $\mathbf{c}$ to generate word $y$. Allowing non-invertibility is a deliberate design choice driven by the fact that the mapping from latent to observed space is often highly complex and unknown. Imposing constraints on it may unnecessarily limit its flexibility. Moreover, in the context of text data, there are two key considerations. First, in natural language processing, the mapping from latent space to observed space often involves a many-to-one relationship. For example, different combinations of emotional concepts can lead to the same sentiment label. In sentiment analysis, the combination of ‘positive sentiment’ and ‘excitement’ might result in an observed outcome such as ‘This is amazing!’ (Miller, 1995). Second, some latent concepts may not be explicitly manifested in the observed sentence. For instance, a speaker’s intent, tone, or implicit connotations may influence how a sentence is constructed, yet remain unobservable in the surface-level text. This phenomenon is particularly prominent in pragmatics and discourse analysis, where unspoken contextual factors significantly shape meaning (Levinson, 1983). By not enforcing invertibility, our model can accommodate these unobservable latent factors, which are essential for robust and piratical identifiability analysis. To this end, we define the approximate invertibility of the mapping $\mathbf{g}$ in Eq. (1) as follows:
@@
We define the degree of approximate invertibility of the mapping $\mathbf{g}$ in Eq. (1) by introducing an error term $\epsilon$, as follows:
$1-p(\mathbf{c}=\mathbf{c}^{*}|\mathbf{x},y)=\epsilon,$
where $10\leq\epsilon<$ 1, and $\mathbf{c}^{*}=\arg\max_{\mathbf{c}}p(\mathbf{c}|\mathbf{x},y)$ represents the dominant point.
@@ 3 Identifiabilty Analysis for the Proposed Latent Variable Model
We now turn to the identifiability analysis of the proposed latent variable model, i.e., whether the latent variables can be uniquely recovered (or recovered up to an equivalence class) from observed data under the given model assumptions. This analysis is conducted within the next-token prediction framework, a widely adopted and empirically significant learning paradigm for training LLMs (Brown et al., 2020; Touvron et al., 2023b; Bi et al., 2024; Zhao et al., 2023).
We begin by introducing the general form of next-token prediction, which serves as the foundation for our analysis. The goal of next-token prediction is to learn a model that predicts the probability distribution over the next token, i.e., $y$, given a context (i.e., input text $\mathbf{x}$). This is achieved by applying the softmax function over the logits produced by the model’s final layer. This process mirrors multinomial logistic regression, where the model approximates the true posterior distribution, i.e., $p(y|\mathbf{x})$, by minimizing the cross-entropy loss. In theory, with sufficient training data, a sufficiently expressive architecture, and proper optimization, the model’s predictions will converge to the true posterior distribution of tokens. Mathematically,
@@
where $y^{\prime}$ denotes a specific value of the observed variable $y$. The function $\mathbf{f}_{\mathbf{x}}(\mathbf{x})$ maps input $\mathbf{x}$ to a representation space, and $\mathbf{f}_{y}(y)$ corresponds to the weights in the model’s final layer. In the following, we will show how the representations $\mathbf{f}_{\mathbf{x}}(\mathbf{x})$ link to the true posterior distribution $p(\mathbf{c}|\mathbf{x})$ in the proposed latent variable model in Eq. (1).
By applying Bayes rule to the proposed latent variable model (Eq. (1)) the true posteriors of the latent variables $\mathbf{c}$ given input $\mathbf{x}$ can be written as:
$p({y}|\mathbf{x})=\sum_{\mathbf{c}}p(y|\mathbf{c})p(\mathbf{c}|\mathbf{x}).$ (3)
Assume that the representation extractor $\mathbf{f}_{\mathbf{x}}$ has a universal approximation capacity, and that the amount of data is infinite, so that next-token prediction converges to the optimal classifier. Then, we will have equality between the posterior Eq. (2) and the true posterior in Eq. (3). Taking the logarithm of the right hand sides of both of Eqs. (2) and (3), we obtain:
$\mathbf{f}_{\mathbf{x}}(\mathbf{x})^{T}\mathbf{f}_{y}(y)-\log Z(\mathbf{x})=\log\sum_{\mathbf{c}}p(y|\mathbf{c})p(\mathbf{c}|\mathbf{x}),$ (4)
where $Z(\mathbf{x})=\sum_{y^{\prime}}{\exp({\mathbf{f}_{\mathbf{x}}(\mathbf{x})^{T}\mathbf{f}_{y}(y^{\prime}))}}$. Equation (4) establishes an initial connection between the inference model (on the left) and the generative model (on the right). Intuitively, the first term on the left is a matrix-vector multiplication, while the right-hand side involves a logarithm of a summation over latent variables.
@@
Without additional assumptions, it is challenging to derive results beyond Eq. (4). To address this, we introduce the following assumption, referred to as the diversity condition. Formally, we assume that: there exist $\ell+1$ distinct values of $y$, where $\ell$ is the number of all possible values that $\mathbf{c}$ can take, i.e., $y_{0}$,…, $y_{k}$, …, $y_{\ell}$, such that the matrix $\mathbf{L}=\big{(}[p(\mathbf{c}=\mathbf{c}_{i}|y=y_{1})]_{\mathbf{c}_{i}}-[p(\mathbf{c}=\mathbf{c}_{i}|y=y_{0})]_{\mathbf{c}_{i}},...,[p(\mathbf{c}=\mathbf{c}_{i}|y=y_{\ell})]_{\mathbf{c}_{i}}-[p(\mathbf{c}=\mathbf{c}_{i}|y=y_{0})]_{\mathbf{c}_{i}}\big{)}$ of size $\ell\times\ell$ is invertible, where $[p(\mathbf{c}=\mathbf{c}_{i}|y=y_{1})]_{\mathbf{c}_{i}}$ denote the vector whose elements correspond to the probabilities of all possible values of $\mathbf{c}$, conditional on $y=y_{1}$. This assumption was initially developed in the context of nonlinear ICA (independent component analysis) (Hyvarinen & Morioka, 2016; Hyvarinen et al., 2019; Khemakhem et al., 2020). Basically, it requires that the distribution $p(\mathbf{c}|y)$ exhibits sufficiently strong and diverse variation across different values of $y$. Similarly, we also assume that there exist $\ell+1$ values of $y$, so that the matrix $\mathbf{\hat{L}}=\big{(}\mathbf{f}_{y}(y=y_{1})-\mathbf{f}_{y}(y=y_{0}),...,\mathbf{f}_{y}(y=y_{\ell})-\mathbf{f}_{y}(y=y_{0})\big{)}$ of size $\ell\times\ell$ is also invertible. This assumption has likewise been employed in the context of identifiability analysis in inference space for LLMs (Roeder et al., 2021; Marconato et al., 2024). Under this diversity condition, we derive the following approximate identifiability result:
@@
$\mathbf{f}_{x}(\mathbf{x})=\mathbf{A}[\log{p(\mathbf{c}=\mathbf{c}_{i}|\mathbf{x})}]_{\mathbf{c}_{i}}-(\mathbf{\hat{L}}^{T})^{-1}\mathbf{{h}}_{y}+(\mathbf{\hat{L}}^{T})^{-1}\mathbf{b}_{y},$ (5)
where ${\mathbf{h}_{y}}={[h_{y_{1}}-h_{y_{0}},...,h_{y_{\ell}}-h_{y_{0}}]}$ with $h_{y_{k}}[{p(\mathbf{c}=\mathbf{c}_{i}|y=y_{k})}]^{T}_{\mathbf{c}_{i}}[\log{p(\mathbf{c}=\mathbf{c}_{i}|y=y_{k},\mathbf{x})}]_{\mathbf{c}_{i}}$, and ${\mathbf{b}_{y}}={[b(y=y_{1})-b(y=y_{0}),...,b(y=y_{\ell})-b(y=y_{0})]}$ with $b(y=y_{k})=\mathbb{E}_{p(\mathbf{c}|y)}[\log{{p(y|\mathbf{c})}}]$, and $\mathbf{A}=(\mathbf{\hat{L}}^{T})^{-1}\mathbf{L}$. When $\epsilon=0$, $\mathbf{f}_{x}(\mathbf{x})=\mathbf{A}[\log{p(\mathbf{c}=\mathbf{c}_{i}|\mathbf{x})}]_{\mathbf{c}_{i}}$. As $\epsilon\rightarrow 0$, $\mathbf{f}_{x}(\mathbf{x})\approx\mathbf{A}[\log{p(\mathbf{c}=\mathbf{c}_{i}|\mathbf{x})}]_{\mathbf{c}_{i}}$.
Differences from Previous Works
Our identifiability result differs from previous studies (Roeder et al., 2021; Marconato et al., 2024; Jiang et al., 2024; Rajendran et al., 2024) in several key aspects. First, our result establishes a novel connection by linking the representations learned through next-token prediction directly to the true latent variables of the generative model. This direct relationship contrasts with prior works (Roeder et al., 2021; Marconato et al., 2024), which primarily compare representations between inference models, leaving the connection to generative latent variables unexplored. Second, our approach relies on relaxed assumptions. Specifically, we assume both latent variables and observed variables to be discrete, aligning more closely with modeling tasks such as natural language, where words are inherently discrete. In contrast, Rajendran et al. (2024) assume continuous latent and observed variables, which may not be appropriate for word-based models. Further, we do not enforce an invertible mapping from latent variables to observed variables, reflecting realistic scenarios in which observed data cannot fully determine the entire latent space. This flexibility is crucial for capturing the complex and often unknown relationships between latent and observed variables, an aspect overlooked in prior works (Rajendran et al., 2024; Jiang et al., 2024), which enforce invertibility.
Further Discussion
Identifiability analysis is a fundamental challenge, especially in the causal representation learning community. While this community often assumes that the mapping from latent space to observed space is invertible (Brehmer et al., 2022; Von Kügelgen et al., 2021; Massidda et al., 2023; von Kügelgen et al., 2023; Ahuja et al., 2023; Seigal et al., 2022; Shen et al., 2022; Liu et al., 2022), such an assumption may, in some real-world applications, oversimplify the inherently complex relationships between latent and observed variables. For example, in financial data, different combinations of latent factors (e.g., interest rates, geopolitical events) can lead to the same outcomes, making the mapping non-invertible. In healthcare, varied genetic and environmental factors can produce identical disease outcomes, while in recommendation systems, multiple latent features can result in the same interaction, all illustrating the non-invertibility of these mappings. These examples illustrate the non-invertibility of such mappings. We hope this work will inspire further research aimed at overcoming the limitations of invertibility assumptions in causal representation learning.

### v3

@@ Abstract
The remarkable achievements of large language models (LLMs) have led many to conclude that they exhibit a form of intelligence. This is as opposed to explanations of their capabilities based on their ability to perform relatively simple manipulations of vast volumes of data. To illuminate the distinction between these explanations, we introduce a novel generative model that generates tokens on the basis of human-interpretable concepts represented as latent discrete variables. Under mild conditions—even when the mapping from the latent space to the observed space is non-invertible—we establish an identifiability result: the representations learned by LLMs through next-token prediction can be approximately modeled as the logarithm of the posterior probabilities of these latent discrete concepts given input context, up to an invertible linear transformation. This theoretical finding not only provides evidence that LLMs capture underlying generative factors, but also provide a unified prospective for understanding of the linear representation hypothesis. Taking this a step further, our finding motivates a reliable evaluation of sparse autoencoders by treating the performance of supervised concept extractors as an upper bound. Pushing this idea even further, it inspires a structural variant that enforces dependence among latent concepts in addition to promoting sparsity. Empirically, we validate our theoretical results through evaluations on both simulation data and the Pythia, Llama, and DeepSeek model families, and demonstrate the effectiveness of our structured sparse autoencoder.
@@ 1 Introduction
Large language models (LLMs) are trained on extensive datasets, primarily sourced from the Internet, enabling them to excel in a wide range of downstream tasks, such as language translation, text summarization, and question answering [78, 10, 58]. This remarkable success has ignited a heated debate surrounding the question of whether their performance is primarily achieved through simple operations applied to vast amounts of memorized data, or is a sign of genuine “intelligence”. Intelligence in this context would imply that the LLM had learned a model of the underlying generative factors that gave rise the data.
Recent empirical evidence has made it increasingly clear that the representations learned by LLMs encapsulate latent concepts, such as sentiment [72] or writing style [49], which align with human-interpretable abstractions [1, 50, 65]. Despite this, LLMs are often viewed as mere memorization systems, with their intelligence questioned due to the fundamental simplicity of next-token prediction. Critics argue that this seemingly trivial approach cannot explain the emergence of advanced AI capabilities, much less form the foundation of artificial general intelligence (AGI).
@@
Figure 1: An overview of the main contributions of this work. On the left, we illustrate the proposed latent variable model that represents concepts as latent variables $\mathbf{c}$, which are used to generate both the input $\mathbf{x}$ and output $y$ within a next-token prediction framework. Leveraging Bayes’ rule, the next-token prediction framework, and the diversity condition, we establish an identifiability result: the representations learned by LLMs approximately correspond to a linear transformation of the logarithm of the posterior distribution of latent variables conditioned on input tokens, i.e., $\mathbf{f}_{\mathbf{x}}(\mathbf{x})=\mathbf{A}[\log{p(\mathbf{c}=\mathbf{c}_{i}|\mathbf{x})}]_{\mathbf{c}_{i}}+\mathbf{b}+o({\epsilon})$, where $\mathbf{b}$ is a constant, and $o(\epsilon)$ represents a term that grows asymptotically smaller than $\epsilon$ as $\epsilon\to 0$. This identifiability result provide a support for the linear representation hypothesis, and sparse autoencoder.
@@
Empirical studies suggest that LLM representations (also known as activations) capture human-interpretable concepts in latent space [72, 1, 50, 65]. To analyze whether and how LLMs learn this through next-token prediction, we introduce a latent variable model, as shown on the left in Figure 1. We model latent concepts as variables connected by arbitrary causal relationships that give rise to observed variables, capturing the underlying generative process between concepts and observations. Although efforts have been taken to model the concepts as continuous linear vectors [63] or explore the empirically recognized linear geometry corresponding to the concepts [59], we aim to develop a model enabling flexibility in capturing the data generative process and theoretical identifiability analysis. We explicitly model all variables, both latent and observed, as discrete, aligning with the inherently discrete nature of language and human-interpretable semantic representations of the world with discreteness. In addition, we do not impose invertibility on the mapping from latent to observed space, as real-world data often involves complex relationships that strict invertibility constraints cannot capture. Building on this model, we develop a theoretical framework to assess whether next-token prediction can learn interpretable concepts. Through identifiability analysis, we show that, under mild conditions, representations learned by next-token prediction approximately correspond to a linear transformation of the logarithm of the posterior distribution of latent variables conditioned on input tokens. This connection offers theoretical insights into LLM mechanisms, highlighting their potential to model aspects of human intelligence. Our linear identifiability result also offers a unified perspective supporting the linear representation hypothesis, which empirically suggests that concepts in LLMs are often encoded linearly [54, 60, 5, 23, 16, 69, 57, 56, 59, 43]. Further, our linear identifiability result motivates a reexamination of sparse autoencoders (SAEs), which model LLM representations as sparse linear combinations of human-interpretable concepts [32, 13]. We propose a new evaluation framework for SAEs, treating supervised concept extraction methods, such as linear probing, as an upper bound. More importantly, we introduce a novel SAE variant, structured SAE, that incorporates structured regularization to account for dependencies among latent concepts, in addition to enforcing sparsity. Our main contributions are summarized as follows:
@@
  • In Section 5, we introduce a novel approach for evaluating SAEs by treating the performance of supervised concept extractors, e.g., linear probing, as an upper bound. Additionally, we propose structured SAEs, a variant that incorporate structured regularization.
  • In Section 6, we present experiments on both simulation data, and real data across the Pythia [8], Llama [70, 18], and DeepSeek-R1 [30] models families, with results consistent with our findings. We also demonstrate the proposed structured SAEs in term of the proposed evaluating method.
@@ 2 A Latent Variable Modal for Text Data
$[\mathbf{x},y]=\mathbf{g}(\mathbf{c}),\quad\text{or probabilistically},\quad p(\mathbf{x},y)=\sum\nolimits_{\mathbf{c}}p(\mathbf{x}|\mathbf{c})p(y|\mathbf{c})p(\mathbf{c}),$ (1)
where latent variables $\mathbf{c}$ model latent human-interpretable concepts, and observed variables $\mathbf{x}$ and $y$ model words generated by the deterministic mapping $\mathbf{g}$ on $\mathbf{c}$. We here distinguish observed variables into $\mathbf{x}$ and $y$ to align the input, i.e., context, and output token in the next-token prediction inference framework, respectively.
Our goal is to impose realistic assumptions on the proposed latent variable model, ensuring the model’s flexibility to closely approximate real language data. To this end, we do not impose any specific graph structure over latent variables. For example, we allow for any directed acyclic graph structure over latent variables from the perspective of causal representation learning. Most importantly, we emphasize the following factors.
@@
Existing works predominantly assume that both latent and observed variables for modeling language data are continuous [63, 77], primarily to simplify identifiability analysis. However, this assumption overlooks the inherently discrete nature of text. In contrast, we assume throughout this work that all variables, both latent variables $\mathbf{c}$ and observed $\mathbf{x}$ and $y$, are discrete. This assumption is particularly natural for observed variables $\mathbf{x}$ and $y$, given the inherently discrete structure of words and categories in text. Likewise, latent variables, which represent underlying concepts, are inherently well-suited to discrete modeling, as they often correspond to categorical distinctions or finite sets of semantic properties that naturally arise in text data. For instance, in a topic modeling scenario, latent variables can represent distinct topics, such as “sports”, “politics”, or “technology”, each of which forms a discrete category. This discrete structure aligns with the way humans typically interpret and classify information in text, making the assumption not only intuitive but also practical for real-world applications [9, 36, 35].
@@
Unlike many existing approaches [35, 63], we do not require the mapping $\mathbf{g}$ from latent to observed variables to be invertible. Allowing non-invertibility is a deliberate design choice driven by the fact that the mapping from latent to observed space is often highly complex and unknown. Imposing constraints on it may unnecessarily limit its flexibility. Moreover, in the context of text data, there are two key considerations. First, in natural language processing, the mapping from latent space to observed space often involves a many-to-one relationship. For example, different combinations of emotional concepts can lead to the same sentiment label. In sentiment analysis, the combination of ‘positive sentiment’ and ‘excitement’ might result in an observed outcome such as ‘This is amazing!’ [55]. Second, some latent concepts may not be explicitly manifested in the observed sentence. For instance, a speaker’s intent, tone, or implicit connotations may influence how a sentence is constructed, yet remain unobservable in the surface-level text. This phenomenon is particularly prominent in pragmatics and discourse analysis, where unspoken contextual factors significantly shape meaning [42]. By not enforcing invertibility, our model can accommodate these unobservable latent factors, which are essential for robust and piratical identifiability analysis. To this end, we define the approximate invertibility of the mapping $\mathbf{g}$ in Eq. (1) as follows:
@@
We define the degree of approximate invertibility of the mapping $\mathbf{g}$ in Eq. (1) by introducing an error term $\epsilon$, as follows: $1-p(\mathbf{c}=\mathbf{c}^{*}|\mathbf{x},y)=\epsilon$, where $10\leq\epsilon<$ 1, and $\mathbf{c}^{*}=\arg\max_{\mathbf{c}}p(\mathbf{c}|\mathbf{x},y)$ represents the dominant mode of the posterior.
@@ 3 Identifiabilty Analysis for the Proposed Latent Variable Model
We now turn to the identifiability analysis of the proposed latent variable model, i.e., whether the latent variables $\mathbf{c}$ can be uniquely recovered (up to an equivalence class) from observed data $\mathbf{x}$ and ${y}$ under additional assumptions. We conduct this analysis within the next-token prediction framework, a widely adopted and empirically significant paradigm for training LLMs [14, 71, 7, 79].
We begin by introducing the general form of next-token prediction, which serves as the foundation for our analysis. The goal of next-token prediction is to learn a model that predicts the conditional distribution over the next token $p(y|\mathbf{x})$, which can be achieved by applying the softmax function over the logits produced by the model’s final layer. This process mirrors multinomial logistic regression, where the model approximates the true conditional distribution $p(y|\mathbf{x})$ by minimizing the cross-entropy loss. In theory, with sufficient training data, a sufficiently expressive architecture, and proper optimization, the model’s predictions will converge to the true conditional distribution. Mathematically,
@@
where $y^{\prime}$ denotes a specific value of the observed variable $y$. The function $\mathbf{f}_{\mathbf{x}}(\mathbf{x})$ maps input $\mathbf{x}$ to a representation space, and $\mathbf{f}_{y}(y)$ corresponds to the weights in the model’s final layer, i.e., look-up table.
On the other hand, the true $p({y}|\mathbf{x})$ can be derived from Eq. (1) using Bayes’ rule:
$p({y}|\mathbf{x})=\sum\nolimits_{\mathbf{c}}p(y|\mathbf{c})p(\mathbf{c}|\mathbf{x}).$ (3)
By the expression for $p({y}|\mathbf{x})$ in both Eq. (2) and Eq. (3), we can align the right-hand sides of these two equations. Taking the logarithm of both sides then yields:
$\mathbf{f}_{\mathbf{x}}(\mathbf{x})^{T}\mathbf{f}_{y}(y)-\log\big{(}\sum\nolimits_{y^{\prime}}{\exp({\mathbf{f}_{\mathbf{x}}(\mathbf{x})^{T}\mathbf{f}_{y}(y^{\prime}))}}\big{)}=\log\sum\nolimits_{\mathbf{c}}p(y|\mathbf{c})p(\mathbf{c}|\mathbf{x}).$ (4)
We now, as shown in Eq. (4), establish a initial connection between the inference model implemented by LLMs (left-hand side) and the generative model (right-hand side): the left involves representations $\mathbf{f}_{\mathbf{x}}$ learned by LLMs, while the right involves the true posterior $p(\mathbf{c}|\mathbf{x})$.
@@
To further investigate the relation between the representations $\mathbf{f}_{\mathbf{x}}$ and the true posterior $p(\mathbf{c}|\mathbf{x})$ in Eq. (4), we introduce the following assumption, referred to as the diversity condition. Formally, we assume that: there exist $\ell+1$ distinct values of $y$, i.e., $y_{0}$,…, $y_{k}$, …, $y_{\ell}$, such that the matrix $\mathbf{L}=\big{(}[p(\mathbf{c}=\mathbf{c}_{i}|y=y_{1})-p(\mathbf{c}=\mathbf{c}_{i}|y=y_{0})]_{\mathbf{c}_{i}},...,[p(\mathbf{c}=\mathbf{c}_{i}|y=y_{\ell})-p(\mathbf{c}=\mathbf{c}_{i}|y=y_{0})]_{\mathbf{c}_{i}}\big{)}$ of size $\ell\times\ell$ is invertible, where $\ell$ is the number of all possible values that $\mathbf{c}$ can take [^1] For example, if $\mathbf{c}=[c^{1},c^{2}]$, and $c^{1}\in\{0,1\},c^{2}\in\{0,1,2\}$, then $\mathbf{c}$ can take $2\times 3$ values., the notation $[\cdot]_{\cdot}$, throughout the paper, denotes the vector indexed by possible values of latent variables. When referring to the full latent variables $\mathbf{c}$, we write $[\cdot]_{\mathbf{c}_{i}}$, when referring to a single concept $c^{i}$, we write $[\cdot]_{c^{i}}$. This assumption was initially developed in the context of nonlinear independent component analysis [33, 34, 39]. Basically, it requires that the distribution $p(\mathbf{c}|y)$ exhibits sufficiently strong and diverse variation across different values of $y$. Similarly, we also assume that there exist $\ell+1$ values of $y$, so that the matrix $\mathbf{\hat{L}}=\big{(}\mathbf{f}_{y}(y=y_{1})-\mathbf{f}_{y}(y=y_{0}),...,\mathbf{f}_{y}(y=y_{\ell})-\mathbf{f}_{y}(y=y_{0})\big{)}$ of size $\ell\times\ell$ is also invertible. This assumption has been employed in the context of identifiability analysis in inference space for LLMs [64, 51]. Under this diversity condition, we derive the following identifiability result:
@@
$\mathbf{f}_{x}(\mathbf{x})=\mathbf{A}[\log{p(\mathbf{c}=\mathbf{c}_{i}|\mathbf{x})}]_{\mathbf{c}_{i}}+\mathbf{b}-(\mathbf{\hat{L}}^{T})^{-1}\mathbf{{h}}_{y},$ (5)
where ${\mathbf{h}_{y}}={[h_{y_{1}}-h_{y_{0}},...,h_{y_{\ell}}-h_{y_{0}}]}$ with $h_{y_{k}}=[{p(\mathbf{c}=\mathbf{c}_{i}|y=y_{k})}]^{T}_{\mathbf{c}_{i}}[\log{p(\mathbf{c}=\mathbf{c}_{i}|y=y_{k},\mathbf{x})}]_{\mathbf{c}_{i}}$, and $\mathbf{b}=(\mathbf{\hat{L}}^{T})^{-1}\mathbf{b}_{y}$, with ${\mathbf{b}_{y}}={[b(y=y_{1})-b(y=y_{0}),...,b(y=y_{\ell})-b(y=y_{0})]}$ and $b(y=y_{k})=\mathbb{E}_{p(\mathbf{c}|y=y_{k})}[\log{{p(y=y_{k}|\mathbf{c})}}]$, and $\mathbf{A}=(\mathbf{\hat{L}}^{T})^{-1}\mathbf{L}$. When $\epsilon=0$, $\mathbf{f}_{x}(\mathbf{x})=\mathbf{A}[\log{p(\mathbf{c}=\mathbf{c}_{i}|\mathbf{x})}]_{\mathbf{c}_{i}}+\mathbf{b}$. As $\epsilon\rightarrow 0$, $\mathbf{f}_{x}(\mathbf{x})\approx\mathbf{A}[\log{p(\mathbf{c}=\mathbf{c}_{i}|\mathbf{x})}]_{\mathbf{c}_{i}}+\mathbf{b}$.
Discussion
Identifiability analysis is a fundamental challenge, especially in the causal representation learning community. While this community often assumes that the mapping from latent space to observed space is invertible [12, 74, 53, 75, 3, 67, 68, 44], such an assumption may, in some real-world applications, oversimplify the inherently complex relationships between latent and observed variables. For example, in financial data, different combinations of latent factors (e.g., interest rates, geopolitical events) may lead to the same outcomes, making the mapping non-invertible. In healthcare, varied genetic and environmental factors may produce identical disease outcomes. These examples illustrate the non-invertibility of such mappings. We hope this work will inspire further research aimed at overcoming the limitations of invertibility assumptions in causal representation learning.